# AMP4EC: Adaptive Model Partitioning Framework for Efficient Deep Learning Inference in Edge Computing Environments thanks:

### v1

@@ I Introduction
Edge computing has emerged as a transformative paradigm in the age of the Internet of Things (IoT) and artificial intelligence (AI), facilitating data processing closer to data sources such as sensors and user devices. By reducing latency and alleviating network bandwidth demands, edge computing enables real-time decision-making and enhances overall system efficiency [1]. However, executing computationally intensive deep learning models within resource-constrained edge environments pose significant challenges [2]. Optimizing limited computational resources while maintaining inference performance remains a critical area of research [3, 4, 5, 6, 7].
Deep neural network (DNN) model partitioning has been proposed as a promising solution to address these challenges. By leveraging the inherently layered structure of DNNs, models can be distributed across various edge computing components, allowing for more efficient resource utilization. For example, Bhattacharya et al. [8] introduced a partial inference scheme that partitions neural networks layer by layer, enabling selective layer processing based on system constraints such as latency and storage. Similarly, Bhardwaj et al. [9] developed the Network of Neural Networks (NoNN) framework, which partitions a teacher model into disjoint, highly compressed student networks, thereby reducing memory and communication overhead while preserving accuracy. Despite their effectiveness, these approaches rely on pre-partitioned knowledge and lack real-time resource monitoring and adaptation mechanisms, limiting their flexibility in dynamic IoT scenarios. For instance, they do not account for resource fluctuations, such as:
@@
  • Device offline: If a device goes offline, the framework lacks the ability to redistribute the computational workload across the remaining devices, which may lead to service degradation or interruptions.
To address these challenges, this paper proposes AMP4EC, an Adaptive Model Partitioning framework designed for efficient deep learning inference in Edge Computing environments. AMP4EC introduces a novel adaptive scheduling mechanism that optimizes task distribution by leveraging real-time resource monitoring and historical performance data. The framework dynamically partitions deep learning models into sub-models using a resource-aware model partitioning algorithm, ensuring alignment with the computational and memory capabilities of heterogeneous edge devices. By incorporating an optimized weighted scoring mechanism, AMP4EC balances resource utilization, fairness, and task performance during scheduling. Additionally, the framework integrates an efficient model deployment strategy to minimize communication overhead and adapt to dynamic resource constraints. Through comprehensive experiments, AMP4EC demonstrates significant improvements in inference latency, throughput, and resource efficiency, establishing it as a scalable and reliable solution for distributed deep learning inference in resource-constrained edge environments.
Experimental evaluations of AMP4EC demonstrated its scalability, adaptability, and efficiency in resource-constrained edge environments. The framework achieved up to a 78% reduction in inference latency and a 415% improvement in throughput compared to a monolithic baseline, while maintaining minimal overhead in scheduling (10ms) and resource monitoring (less than 1% CPU utilization). Its adaptive task scheduling and model partitioning strategies effectively balanced workloads across heterogeneous edge nodes, achieving linear performance scaling with up to three nodes. Additionally, AMP4EC exhibited robust performance under varying resource profiles, demonstrating efficient inference times in high and medium configurations (234.56ms and 389.27ms, respectively) and adaptability to low-resource environments (583.91ms). These results underscore AMP4EC’s potential as a scalable and efficient solution for distributed deep learning inference in dynamic and resource-constrained edge computing scenarios.
@@ II Background and Related Work
Over the recent decade, the emergence of the new generation of IoT devices and edge computing in a resource-constrained setting have necessitated novel techniques to allow efficient training of deep learning models and their deployment at the edge of networks [10, 11, 12, 13, 14]. Traditional centralized learning systems, however effective in high-resource environments, are heavily limited in scenarios requiring low latency, data privacy, and scalability. In the effort to tackle these challenges, researches both on model optimization methods and hardware innovations have been conducted. One such approach, Distributed Deep Learning (DDL) systems, have gained great traction as an enabling paradigm. This approach distributes deep learning tasks and computational workloads over multiple computing nodes, including edge servers and centralized server, and it has achieved low-latency high-efficiency training and inference [5, 6, 7].
@@
Various approaches have long been undertaken to enhance DDL performance across distributed environments. Lane et al. [15] made an early attempt to assess the feasibility of deep learning on embedded devices with a prototype of low-power Deep Neural Network (DNN) inference engine on a mobile device in 2015. Various techniques have been proposed to address these complexities of distributed training and inference. Federated Learning (FL) stands out as an attractive framework known for its collaborative and privacy-preserving attribute. FL supports the training over client-sided devices in a decentralized manner followed by central aggregating of the updates. Geyer et al. [16] presented a FL framework that integrates a Differential Privacy (DP) mechanism. While effective, FL often suffers from issues such as high communication overhead. An alternative approach is Split Learning (SL), which partitions training tasks between clients and servers, but at the cost of increased communication latency. Since then, more advanced derivatives such as Parallel Split Learning (PSL) and Split Federated Learning (SFL) have been proposed to overcome these limitations. For instance, Lin et al. [17] advocated a PSL method which parallelizes client-side model training via last-layer gradient aggregation.
Although they have been known for workload distributiveness and scalability, DDL systems often depend on fixed computation and partitioning strategies and do not effectively exploit and adapt the resource heterogeneity at the edge. The challenges include addressing heterogeneity in computational power, network bandwidth, and storage while ensuring effective coordination and task allocation. On this front, model partitioning has emerged as a critical technique, allowing DNN to be divided into segments that can be dynamically executed across devices, adapting to resource constraints and enhancing performance in distributed environments.
@@
Depending on the research focuses and corresponding challenges, model partitioning techniques can be classified in various ways [18]. We outlined two dimensions of partitioning in the context of our proposal—training phase partitioning versus inference phase partitioning, and static partitioning versus resource-aware partitioning. The scope of this work focuses on inference phase and resource-aware partitioning, where the aim is to optimize real-time inference performance in resource-constrained environments. In static partitioning, splits of DNNs are pre-determined and remain unchanged throughout the execution. While this simplifies implementation, it lacks adaptability to changes in resource availability. For instance, Dey et al. [19] proposed a layer-wise partitioning strategy that divides the model into segments for selective processing based on latency and storage constraints. Resource-aware partitioning, in contrast, adjusts model splitting strategies to the specific constraints of edge devices, dynamically monitoring memory, CPU, network latency and system stability. Kim et al. [20]introduced a memory-aware DNN partitioning framework that enhances inference efficiency by dynamically assigning layers based on device profiles. Similarly, Fan et al. [21]developed an energy-aware partitioning method to minimize energy consumption in edge-to-cloud distributed systems. However, we found that existing works explore resource-aware strategies but few monitors and optimizes the environment of multiple resource dimensions simultaneously in critical scenarios where edge environments undergo dynamic changes, such as the integration of new devices or the failure of existing ones. Fluctuating environment in real time can cause inefficiency or interruption of inference-grounding, highlighting the need for adaptive frameworks capable of redistributing workloads in real time.
@@ III-B Model Partitioner (B)
The Model Partitioner is responsible for analyzing a deep learning model, estimating its computational and resource requirements, and dividing it into partitions for distributed execution. This partitioning ensures balanced workloads across edge devices while minimizing communication overhead.
The main functionalities of the Model Partitioner are as follows:
@@ III-B3 Partition Boundaries (B3)
Based on the cumulative computational cost, the model is divided into a specified number of partitions (num_partitions). Each partition balances the workload across devices by considering both computational cost and communication overhead, with dynamic adjustment based on node capabilities:
@@ IV-A Experimental Setup
A pre-trained model MobileNetV2 [22] was employed for evaluation, leveraging its efficient architecture and compatibility with a wide range of edge servers. The evaluation was conducted over 100 inference iterations for each configuration to ensure statistical robustness and repeatability of the results. Resource profiles were categorized into three configurations: High (1 CPU, 1GB), Medium (0.6 CPU, 512MB), and Low (0.4 CPU, 512MB), representing varying computational and memory constraints. Key metrics collected from the experiment include Inference Latency, Throughput (req/s), Communication Overhead (ms), CPU Usage (percent), Memory Usage (MB), Network Bandwidth, Stability Score, Scheduling Overhead (ms). This setup provided an isolated and consistent environment for running the experiments providing comprehensive insights into the performance under varying resource constraints.
@@ IV-B Comparative Performance Analysis
To evaluate the effectiveness of AMP4EC, we conducted comprehensive experiments across three system configurations: (1) a monolithic baseline deployment running on a single node, (2) the basic AMP4EC enhanced with Node Selection Algorithm (NSA) for intelligent task distribution but without caching, and (3) AMP4EC enhanced with NSA and with caching (AMP4EC+Cache). The experiments utilized MobileNetV2 as the test model, with each configuration processing identical batches of 32 inference requests. The monolithic baseline was deployed on a single container with 2 cores and 2GB memory, while distributed configurations leveraged a heterogeneous cluster of Docker containers simulating edge devices with varying computational capabilities (1.0 core / 1GB, 0.6 core / 512MB and 0.4 core / 512MB). Resource constraints were enforced using cgroups, and performance metrics were sampled at 1-second intervals using Docker stats API. Each experiment included a 30-second warm-up period followed by a 5-minute evaluation phase, with measurements repeated three times to ensure statistical significance.

### v2

@@ I Introduction
Edge computing has emerged as a transformative paradigm in the age of the Internet of Things (IoT) and artificial intelligence (AI), facilitating data processing closer to data sources such as sensors and user devices. By reducing latency and alleviating network bandwidth demands, edge computing enables real-time decision-making and enhances overall system efficiency [1]. However, executing computationally intensive deep learning models within resource-constrained edge environments pose significant challenges [2]. Optimizing limited computational resources while maintaining inference performance remains a critical area of research [3, 4, 5, 6, 7, 8].
Deep neural network (DNN) model partitioning has been proposed as a promising solution to address these challenges. By leveraging the inherently layered structure of DNNs, models can be distributed across various edge computing components, allowing for more efficient resource utilization. Existing approaches, such as layer-wise model partitioning [9] and the Network of Neural Networks (NoNN) framework [10], demonstrate effective resource utilization and reduced memory and communication overhead. However, these methods rely on pre-partitioned knowledge and lack real-time resource monitoring and adaptation mechanisms, limiting their flexibility in dynamic IoT scenarios. Specifically, they fail to account for resource fluctuations, such as:
@@
  • Device offline: If a device goes offline, the framework lacks the mechanisms for redistributing workloads, potentially leading to system inefficiencies or inference failures.
To address these challenges, this paper proposes AMP4EC, an Adaptive Model Partitioning framework designed for efficient deep learning inference in Edge Computing environments. AMP4EC introduces a novel adaptive scheduling mechanism that optimizes task distribution by leveraging real-time resource monitoring and historical performance data. Through comprehensive experiments, AMP4EC demonstrates significant improvements in inference latency, throughput, and resource efficiency, establishing it as a scalable and reliable solution for distributed deep learning inference in resource-constrained edge environments.
Experimental evaluations of AMP4EC demonstrated its scalability, adaptability, and efficiency in resource-constrained edge environments. The framework achieved up to a 78% reduction in inference latency and a 415% improvement in throughput compared to a monolithic baseline, while maintaining minimal scheduling overhead (10ms) and negligible CPU utilization (<1%). Its adaptive task scheduling and model partitioning strategies effectively balanced workloads across heterogeneous edge nodes, achieving linear performance scaling with up to three nodes. Additionally, AMP4EC exhibited robust performance under varying resource profiles, demonstrating efficient inference times in high and medium configurations (234.56ms and 389.27ms, respectively) and adaptability to low-resource environments (583.91ms). These results underscore AMP4EC’s potential as a scalable and efficient solution for distributed deep learning inference in dynamic and resource-constrained edge computing scenarios.
@@ II Background and Related Work
Over the recent decade, the emergence of the new generation of IoT devices and edge computing in a resource-constrained setting have necessitated novel techniques to allow efficient training of deep learning models and their deployment at the edge of networks [1, 11, 12, 13, 14, 15]. Traditional centralized learning systems, however effective in high-resource environments, struggles with low latency, data privacy, and scalability requirements. Distributed Deep Learning (DDL) systems, have gained great traction as an enabling paradigm, distributing deep learning tasks and computational workloads across edge servers and centralized server to achieve low-latency high-efficiency training and inference [5, 6, 7].
@@
Various approaches have long been undertaken to enhance DDL performance across distributed environments. Lane et al. [16] made an early attempt to assess the feasibility of deep learning on embedded devices with a prototype of low-power Deep Neural Network (DNN) inference engine on a mobile device in 2015, leading to frameworks like Federate Learning(FL) and Split Learning(SL). FL enables decentralized training with privacy preservation but suffers from high communication overhead [17], while SL partitions tasks between clients and servers at the cost of increased latency. Advanced derivatives such as Parallel Split Learning (PSL) [18] and Split Federated Learning (SFL) have been proposed to address some of these inefficiencies but remain constrained by static computation strategies.
@@
Model partitioning has emerged as a solution to improve DDL adaptability, allowing deep neural networks(DNNs) to be segmented and executed dynamically across devices. Depending on the research focuses and corresponding challenges, model partitioning techniques can be classified in various ways [19]. We outlined two dimensions of partitioning: training versus inference phase partitioning, and static versus resource-aware partitioning. The scope of this work focuses on inference phase and resource-aware partitioning, where the aim is to optimize real-time inference performance in resource-constrained environments. Unlike static approaches such as Dey et al. [20], whci predefine model splits based on latency ans storage, resource-aware techniques optimize allocations in real time. Kim et al. [21]introduced a memory-aware DNN partitioning framework that enhances inference efficiency by dynamically assigning layers based on device profiles. Similarly, Fan et al. [22]developed an energy-aware partitioning method to minimize energy consumption in edge-to-cloud distributed systems. However, existing solutions explore resource-aware strategies but fail to monitor and optimize multiple resource dimensions simultaneously, limiting their adaptability to dynamic edge environments such as the integration of new devices or the failure of existing ones. Fluctuating environment in real time can cause inefficiency or interruption of inference-grounding, highlighting the need for adaptive frameworks capable of redistributing workloads in real time.
@@ III-B Model Partitioner (B)
The Model Partitioner is responsible for analyzing a deep learning model, estimating its computational and resource requirements, and dividing it into partitions for distributed execution. This partitioning ensures balanced workloads across edge devices while minimizing communication overhead. Unlike conventional quantization techniques (e.g., INT8 and FP16 quantization), which focus on reducing model size, our approach employs a Resource-Aware Layerwise Optimization Strategy (RALOS), which better adapts to the heterogeneous resource constraints of edge devices while maintaining the original model accuracy. This method simultaneously considers computation, memory, and communication in the optimization process, employing a weighted scoring mechanism in partition optimization, effectively reducing end-to-end latency by 27.5%.
The main functionalities of the Model Partitioner are:
@@ III-B3 Partition Boundaries (B3)
Based on the cumulative computational cost, the model is divided into a specified number of partitions (num_partitions). Our optimization workflow comprises three key steps:
  • Node Capability Analysis: Calculate each node’s computational capability score:
    $S_{i}=w_{\text{cpu}}\cdot c_{i}+w_{\text{mem}}\cdot M_{I}$ (3)
  • Capability-based Layer Allocation: Distribute model layers according to the ratio:
    $P_{i}=\frac{S_{i}}{\sum_{j}S_{j}}$ (4)
  • Load Balancing Optimization: Iteratively adjusted to reduce communication overhead, measure partition balance using:
    $L=\frac{1}{n}\sum_{i=1}^{n}\left|L_{i}-L_{\text{avg}}\right|$ (5)
Each partition balances the workload across devices by considering both computational cost and communication overhead, with dynamic adjustment based on node capabilities:
@@ III-C Task Scheduler (C)
Algorithm 1 has a computational complexity of $\mathcal{O}(m\cdot n)$, where $m$ is the number of tasks and $n$ is the number of available edge nodes. In practice, each node selection requires calculating scores for $\mathcal{O}(n)$ nodes, totaling $\mathcal{O}(m\cdot n)$ operations for $m$ tasks. As the edge network scales, this complexity has important implications for resource-constrained IoT and edge devices:
  1. 1.
    In small deployments ($n<10$), scheduling overhead is negligible, typically under 1 ms.
  2. 2.
    In medium-scale deployments ($5010\leq n<$ 50), the overhead grows linearly but remains acceptable (1–10 ms).
  3. 3.
    In large-scale deployments ($n\geq 50$), a hierarchical scheduling strategy can be employed to optimize the complexity to $\mathcal{O}(m\cdot\log n)$.
[…]
@@ IV-A Experimental Setup
A pre-trained model MobileNetV2 [23] was employed for evaluation, leveraging its efficient architecture and compatibility with a wide range of edge servers. The evaluation was conducted over 100 inference iterations for each configuration to ensure statistical robustness and repeatability of the results. Resource profiles were categorized into three configurations: High, Medium, and Low as in Table  II, representing varying computational and memory constraints. Key metrics collected from the experiment include Inference Latency, Throughput (req/s), Communication Overhead (ms), CPU Usage (percent), Memory Usage (MB), Network Bandwidth, Stability Score, Scheduling Overhead (ms). This setup provided an isolated and consistent environment for running the experiments providing comprehensive insights into the performance under varying resource constraints.
@@ IV-B Comparative Performance Analysis
To evaluate the effectiveness of AMP4EC, we conducted comprehensive experiments across three system configurations: (1) a monolithic baseline deployment running on a single node, (2) the basic AMP4EC enhanced with Node Selection Algorithm (NSA) for intelligent task distribution but without caching, and (3) AMP4EC enhanced with NSA and with caching (AMP4EC+Cache). The experiments utilized MobileNetV2 as the test model, with each configuration processing identical batches of 32 inference requests. The monolithic baseline was deployed on a single container with 2 cores and 2GB memory, while distributed configurations leveraged a heterogeneous cluster of Docker containers simulating edge devices with varying computational capabilities shown in Table II. Resource constraints were enforced using cgroups, and performance metrics were sampled at 1-second intervals using Docker stats API. Each experiment included a 30-second warm-up period followed by a 5-minute evaluation phase, with measurements repeated three times to ensure statistical significance.